A solid cube has its superficialExpansivity as 1.01*10-6 K-1 Find it's cubic expansivity

1
Expert's answer
2022-01-26T15:31:17-0500

Since superficial expansivity is 2 times greater than linear:


"\beta=2\alpha."


Linear and cubic expansivities:


"\gamma=3\alpha."


Rearrangement gives us


"\gamma=\dfrac 32\beta=1.52·10^{-6}\text{ 1/K}."